OpenStudy (anonymous):

i am stuck on the following if x^2-y^2=a and x-y=b, then what does x^2+y^2=

OpenStudy (anonymous):

the suspense is killing me

hero (hero):

:o)

hero (hero):

aw

OpenStudy (dumbcow):

Difference of squares x^2 - y^2 = (x+y)(x-y) a = b*(x+y) --> x+y = a/b haha sorry got stuck here

OpenStudy (anonymous):

yeah...

hero (hero):

so what does x^2 + y^2 equal again?

myininaya (myininaya):

lol i think

hero (hero):

wait isn't it (a/b)^2

OpenStudy (anonymous):

because \[\frac{a}{b}=(x+y)\] so \[(\frac{a}{b})^2=(x+y)^2=x^2+y^2+2xy\]

OpenStudy (anonymous):

i mean that is why it is not (a/b)^2

myininaya (myininaya):

\[x^2-y^2=a =>x^2=a+y^2\] \[x-y=b=>x=b+y=>x^2=(b+y)^2=b^2+2by+y^2\] \[b^2+2by+y^2=a+y^2=b^2+2by=a=>a-b^2=2by=>y=\frac{a-b^2}{2b}\] but we have \[x^2=a+y^2=>x^2+y^2=a+y^2+y^2=a+2y^2=a+2 \cdot (\frac{a-b^2}{2b})^2\]

OpenStudy (anonymous):

the answer choices are a. 2xy b. -2xy c. (a/b)^2 +2xy d. (a/b)^2-2xy e. (a/b) - xy

OpenStudy (anonymous):

Thanks for everyone's thoughts on this, it is really confusing me

OpenStudy (anonymous):

the answer choices would have been helpful in the begining lol

hero (hero):

lol

OpenStudy (anonymous):

I know, sorry about that

hero (hero):

d

hero (hero):

That's why I meant, I meant d all along

myininaya (myininaya):

\[a+2 \cdot (\frac{a-b^2}{2b})^2=a+2 \cdot \frac{a^2-2ab^2+b^4}{4b^2}\] \[=a+\frac{a^2-2ab^2+b^4}{2b^2}=\frac{2ab^2+a^2-2ab^2+b^4}{2b^2}\] \[=\frac{a^2+b^4}{2b^2}\] wait our answer could have x and y in it ? i didn't know that

hero (hero):

I'm pretty certain the answer is d

OpenStudy (anonymous):

x^2-y^2=a (x+y)(x-y)=a and x-y=b so (x+y)b=a x+y=a/b (x+y)^2=(a/b)^2 x^2+2xy+y^2=(a/b)^2 x^2+y^2-2xy+2xy=(a/b)^2-2xy x^2+y^2=(a/b)^2-2xy

myininaya (myininaya):

my answer is cooler than all of those

myininaya (myininaya):

lol

OpenStudy (anonymous):

yes it is actually

hero (hero):

jahtoday, I like that. Simple and to the point

OpenStudy (anonymous):

Thank you all!!! This great, I have spent hours on this

hero (hero):

myininaya, if I could give you five medals, I would

myininaya (myininaya):

lol

OpenStudy (anonymous):

PLEASE!!! SOME ONE HELP!!! http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e499e390b8b8d00ebe47725

OpenStudy (anonymous):

i love it! answer had a, b, x and y! lol

OpenStudy (anonymous):

we had \[(\frac{a}{b})^2=x^2+y^2+2xy\] and didn't know how to get rid of the 2xy part. turns out you just need to subtract! who knew?

OpenStudy (anonymous):

lol, myininaya's answer is classier imo.

myininaya (myininaya):

thanks joe :)